Taybae:

Lets see if y’all hella smart 😂. Which number is equivalent to 3^(4)÷3^(2)

Midnight97:

It is an problem

Midnight97:

Not an number

toga:

3^(4) ÷ 3^(2) can be simplified as follows: 3^(4) ÷ 3^(2) = 3^(4-2) = 3^2 = 9. Therefore, the number equivalent to 3^(4) ÷ 3^(2) is 9.